# Smooth Tests for Normality in ANOVA

### v3

@@ Appendix A Appendix
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)=\mathrm{E}\left[\pi_{k}(Z)\right]+o_{p}(1),$
as $N\to\infty$, which implies $\hat{\Psi}_{K}^{2}\to_{p}\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}$. Furthermore,
$\sqrt{N}\left(\hat{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right)$
$=$ $\sqrt{N}\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right]^{\top}\bm{\Sigma}_{K}^{-1}\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right]$
@@
$=$ $2a^{\top}\bm{\Sigma}_{K}^{-1}\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)+\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)\right]^{\top}\bm{\Sigma}_{K}^{-1}$
$\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)-\bm{a}_{K}\right)\right].$ (A.11)
@@
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}(\hat{Z}_{ij})-\bm{a}_{K}\right).$
@@
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\pi_{k}(\hat{Z}_{ij})-a_{k}\right)=\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\zeta_{k}(e_{ij})-\mathrm{E}\left[\zeta_{k}(e_{ij})\right]\right)+o_{p}(1),$
@@
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}(\hat{Z}_{ij})-\bm{a}_{K}\right)\to_{d}\mathcal{N}_{K}\left(\bm{0},\bm{\Xi}_{K}\right),$
@@
$\xi_{kl}=$ $\mathrm{E}\left[\zeta_{k}(e)\zeta_{l}(e)\right]$
$=$ $\mathrm{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\left[d_{1k}d_{3l}+d_{1l}d_{3k}\right]+d_{1k}d_{1l}+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]$
$-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{1}{2}\left[d_{1k}d_{2l}+d_{1l}d_{2k}\right]\mathrm{E}\left[e^{3}\right]+\frac{d_{2k}d_{2l}}{4}\left[\mathrm{E}\left[e^{4}\right]-1\right].$
@@
We first examine the convergence of $\hat{\mu}-\mu$ and $\hat{\sigma}^{2}-\sigma^{2}$ as before. For $\hat{\mu}-\mu$, (A.1) still holds under $H_{1L}$, and $\sqrt{N}(\hat{\mu}-\mu)\to_{d}\mathcal{N}(0,\sigma^{2})$ as $N\to\infty$. For $\hat{\sigma}^{2}-\sigma^{2}$, under $H_{1L}$, (A.2) still holds, and $\sqrt{N}(\hat{\sigma}^{2}-\sigma^{2})\to_{d}\mathcal{N}(0,2\sigma^{4})$  [^1] Here under $H_{1L}$, the asymptotic normality of $\sqrt{N}(\hat{\mu}-\mu)$ and $\sqrt{N}(\hat{\sigma}^{2}-\sigma^{2})$ are both derived from CLT for triangular arrays since the distribution of $e_{ij}$ changes over $N$. It can be verified that the conditions of Lindeberg–Feller CLT hold for both random sequences, see, for example, Proposition 2.27 of Van der Vaart, (2000) . since
$\mathrm{E}\left[e^{4}\right]=$ $\int_{-\infty}^{\infty}x^{4}f(x)\mathrm{d}x$
@@
Next, we clarify the relationship of the denoted constants. Recall the definitions of the constants $d_{1k}=\mathrm{E}[\dot{\pi}_{k}(Z)\phi\left(e\right)]=\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))\phi\left(x\right)f(x)\mathrm{d}x$ and $d_{2k}=\mathrm{E}[\dot{\pi}_{k}(Z)\phi\left(e\right)e]=\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))\phi\left(x\right)xf(x)\mathrm{d}x$. Then under $H_{1L}$, as $N\to\infty$, we have
@@
$\mathrm{E}[\pi_{k}(Z)]=$ $\int_{-\infty}^{\infty}\pi_{k}\left(\Phi(x)\right)f(x)\mathrm{d}x$
@@
Now we start from (A.3) again. Under $H_{1L}$, we still have (A.3), (A.4) and (A.5). (A.6) also holds as a result of (A.12) and (A.13). The remainder terms $R_{N1}$ in (A.7) and $R_{N2}$ in (A.8) are still $O_{p}(N^{-1/2})$. Combining (A.3), (A.4), (A.5), (A.6), (A.7), (A.8) and (A.14), we obtain (3.8) in Theorem 3.
It suffices to derive the asymptotic normality of $N^{-1/2}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\hat{Z}_{ij})$. Under $H_{1L}$, straightforward calculation yields that
$\mathrm{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]=$ $\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))\pi_{l}(\Phi(x))f(x)\mathrm{d}x$
@@
$d_{3k}=$ $\mathrm{E}\left[\pi_{k}(Z)e\right]=\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))xf(x)\mathrm{d}x$
@@
$d_{4k}=$ $\mathrm{E}\left[\pi_{k}(Z)e^{2}\right]=\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))x^{2}f(x)\mathrm{d}x$
@@
$\mathrm{E}\left[e^{3}\right]=$ $\int_{-\infty}^{\infty}x^{3}f(x)\mathrm{d}x$
@@
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\hat{Z}_{ij}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\hat{Z}_{ij}\right),\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\hat{Z}_{ij}\right)\right]^{\top}$
$\to_{d}\mathcal{N}_{K}\left(\bm{\Delta}_{K},\bm{\Sigma}_{K}\right),$
@@
$N\hat{\Psi}_{K}^{2}\to_{d}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}\right),$
@@
Proof of Theorem 4, 5 and 6.
The proof is similar to that of Theorems 1, 2, and 3. Note that in the case of different means and same variance, we assume $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. To simplify the procedure and notations, here we just clarify the asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\hat{Z}_{ij})$ without distinction of the null and alternatives. Our goal is to prove that
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\hat{Z}_{ij}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]e_{ij}-\frac{\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]}{2}\left(e_{ij}^{2}-1\right)\right\}$
@@
We first focus on the estimation effects of $\hat{\mu}_{j}-\mu_{j}$ and $\hat{\sigma}^{2}-\sigma^{2}$. For $\hat{\mu}_{j}-\mu_{j}$, by CLT, we have $\sqrt{N_{j}}(\hat{\mu}_{j}-\mu_{j})\to_{d}\mathcal{N}(0,\sigma^{2})$ for each $1\leq j\leq J$ as $\min\{N_{1},\ldots,N_{J}\}\to\infty$. Besides, since
$\mathrm{E}\left[\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{2}\right]=\frac{\sigma^{2}}{N}\sum_{j=1}^{J}\mathrm{E}\left[N_{j}\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{2}\right]=\frac{J}{N}\sigma^{2},$ (A.16)
we have $N^{-1}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{2}=O_{p}(J/N)=o_{p}(N^{-1/2})$ from $J=o(N^{1/2})$. Analogously, by
$\mathrm{E}\left[\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{4}\right]=\frac{1}{N_{j}^{4}}\left[3N_{j}(N_{j}-1)+N_{j}\mathrm{E}\left[e^{4}\right]\right]=\frac{3}{N_{j}^{2}}+\frac{\mathrm{E}\left[e^{4}\right]-3}{N_{j}^{3}},$
@@
$\mathrm{E}\left[\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{4}\right]$
$=$ $\frac{\sigma^{4}}{N}\sum_{j=1}^{J}\mathrm{E}\left[N_{j}\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{4}\right]=\frac{\sigma^{4}}{N}\sum_{j=1}^{J}\left(\frac{3}{N_{j}}+\frac{\mathrm{E}\left[e^{4}\right]-3}{N_{j}^{2}}\right),$ (A.17)
which implies $N^{-1}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{4}=N^{-1}O_{p}(\sum_{j=1}^{J}N_{j}^{-1})=o_{p}(N^{-1})$ since $\sum_{j=1}^{J}N_{j}^{-1}=o_{p}(1)$. For $\hat{\sigma}^{2}-\sigma^{2}$, we have
$\hat{\sigma}^{2}-\sigma^{2}=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\left(Y_{ij}-\hat{\mu}_{j}\right)^{2}-\sigma^{2}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\varepsilon_{ij}^{2}-\sigma^{2}\right)-\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)^{2}$
@@
$\frac{1}{N\sigma}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}$
$=$ $\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}+o_{p}\left(\frac{1}{\sqrt{N}}\right).$ (A.18)
@@
$\frac{1}{N\sigma}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}$
@@
$\frac{1}{N\sigma}\sum_{j=1}^{J}N_{j}\left(\hat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left(\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{ij}$
$+\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}\neq i_{1}}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}$
@@
$\mathrm{E}[A_{N1}]=\frac{J}{N}\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right],$
@@
$\operatorname{Var}[A_{N1}]=\operatorname{Var}\left[\left(\dot{\pi}_{k}(Z)\phi\left(e\right)-\mathrm{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e\right]\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{1}{N_{j}},$
then $A_{N1}=o_{p}(N^{-1/2})$ since $J=o(N^{1/2})$ as $N\to\infty$. For $A_{N2}$, we have $\mathrm{E}[A_{N2}]=0$, and

### v4

@@ A.2 Proof of Theorem 2
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)=\mathbb{E}\left[\pi_{k}(Z)\right]+o_{p}(1),$
as $N\to\infty$, which implies $\widehat{\Psi}_{K}^{2}\overset{p}{\to}\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}$. Furthermore,
$\sqrt{N}\left(\widehat{\Psi}_{K}^{2}-\bm{a}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{a}_{K}\right)$
$=$ $\sqrt{N}\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right]^{\top}\bm{\Sigma}_{K}^{-1}\left[\bm{a}_{K}+\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right]$
@@
$=$ $2a^{\top}\bm{\Sigma}_{K}^{-1}\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right)+\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right)\right]^{\top}\bm{\Sigma}_{K}^{-1}$
$\left[\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)-\bm{a}_{K}\right)\right].$ (A.11)
@@
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}(\widehat{Z}_{ij})-\bm{a}_{K}\right).$
@@
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\pi_{k}(\widehat{Z}_{ij})-a_{k}\right)=\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\zeta_{k}(e_{ij})-\mathbb{E}\left[\zeta_{k}(e_{ij})\right]\right)+o_{p}(1),$
@@
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\bm{\pi}_{K}(\widehat{Z}_{ij})-\bm{a}_{K}\right)\overset{d}{\to}\mathcal{N}_{K}\left(\bm{0},\bm{\Xi}_{K}\right),$
@@
$\xi_{kl}=$ $\mathbb{E}\left[\zeta_{k}(e)\zeta_{l}(e)\right]$
$=$ $\mathbb{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]-a_{k}a_{l}-\left[d_{1k}d_{3l}+d_{1l}d_{3k}\right]+d_{1k}d_{1l}+\frac{1}{2}\left[a_{k}d_{2l}+a_{l}d_{2k}\right]$
$-\frac{1}{2}\left[d_{2k}d_{4l}+d_{2l}d_{4k}\right]+\frac{1}{2}\left[d_{1k}d_{2l}+d_{1l}d_{2k}\right]\mathbb{E}\left[e^{3}\right]+\frac{d_{2k}d_{2l}}{4}\left[\mathbb{E}\left[e^{4}\right]-1\right].$
@@ A.3 Proof of Theorem 3
We first examine the convergence of $\widehat{\mu}-\mu$ and $\widehat{\sigma}^{2}-\sigma^{2}$ as before. For $\widehat{\mu}-\mu$, (A.1) still holds under $H_{1L}$, and $\sqrt{N}(\widehat{\mu}-\mu)\overset{d}{\to}\mathcal{N}(0,\sigma^{2})$ as $N\to\infty$. For $\widehat{\sigma}^{2}-\sigma^{2}$, under $H_{1L}$, (A.2) still holds, and $\sqrt{N}(\widehat{\sigma}^{2}-\sigma^{2})\overset{d}{\to}\mathcal{N}(0,2\sigma^{4})$  [^1] Here under $H_{1L}$, the asymptotic normality of $\sqrt{N}(\widehat{\mu}-\mu)$ and $\sqrt{N}(\widehat{\sigma}^{2}-\sigma^{2})$ are both derived from CLT for triangular arrays since the distribution of $e_{ij}$ changes over $N$. It can be verified that the conditions of Lindeberg–Feller CLT hold for both random sequences, see, for example, Proposition 2.27 of van2000asymptotic . since
$\mathbb{E}\left[e^{4}\right]=$ $\int_{-\infty}^{\infty}x^{4}f(x)\mathrm{d}x$
@@
Next, we clarify the relationship of the denoted constants. Recall the definitions of the constants $d_{1k}=\mathbb{E}[\dot{\pi}_{k}(Z)\phi\left(e\right)]=\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))\phi\left(x\right)f(x)\mathrm{d}x$ and $d_{2k}=\mathbb{E}[\dot{\pi}_{k}(Z)\phi\left(e\right)e]=\int_{-\infty}^{\infty}\dot{\pi}_{k}(\Phi\left(x\right))\phi\left(x\right)xf(x)\mathrm{d}x$. Then under $H_{1L}$, as $N\to\infty$, we have
@@
$\mathbb{E}[\pi_{k}(Z)]=$ $\int_{-\infty}^{\infty}\pi_{k}\left(\Phi(x)\right)f(x)\mathrm{d}x$
@@
Now we start from (A.3) again. Under $H_{1L}$, we still have (A.3), (A.4), and (A.5). Also, (A.6) holds as a result of (A.12) and (A.13). The remainder terms $R_{N1}$ in (A.7) and $R_{N2}$ in (A.8) are still $O_{p}(N^{-1/2})$. Combining (A.3), (A.4), (A.5), (A.6), (A.7), (A.8), and (A.14), we obtain (3.8) in Theorem 3.
It suffices to derive the asymptotic normality of $N^{-1/2}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}(\widehat{Z}_{ij})$. Under $H_{1L}$, straightforward calculation yields that
$\mathbb{E}\left[\pi_{k}(Z)\pi_{l}(Z)\right]=$ $\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))\pi_{l}(\Phi(x))f(x)\mathrm{d}x$
@@
$d_{3k}=$ $\mathbb{E}\left[\pi_{k}(Z)e\right]=\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))xf(x)\mathrm{d}x$
@@
$d_{4k}=$ $\mathbb{E}\left[\pi_{k}(Z)e^{2}\right]=\int_{-\infty}^{\infty}\pi_{k}(\Phi(x))x^{2}f(x)\mathrm{d}x$
@@
$\mathbb{E}\left[e^{3}\right]=$ $\int_{-\infty}^{\infty}x^{3}f(x)\mathrm{d}x$
@@
$\frac{1}{\sqrt{N}}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\bm{\pi}_{K}\left(\widehat{Z}_{ij}\right)$ $=\sqrt{N}\left[\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{1}\left(\widehat{Z}_{ij}\right),\ldots,\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{K}\left(\widehat{Z}_{ij}\right)\right]^{\top}$
$\overset{d}{\to}\mathcal{N}_{K}\left(\bm{\Delta}_{K},\bm{\Sigma}_{K}\right),$
@@
$N\widehat{\Psi}_{K}^{2}\overset{d}{\to}\chi_{K}^{2}\left(\bm{\Delta}_{K}^{\top}\bm{\Sigma}_{K}^{-1}\bm{\Delta}_{K}\right),$
@@
A.4 Proof of Theorem 4, 5 and 6
The proof is similar to that of Theorems 1, 2, and 3. Note that in the case of different means and same variance, we assume $\min\{N_{1},\ldots,N_{J}\}\to\infty$, $J=o(N^{1/2})$ and $\sum_{j=1}^{J}N_{j}^{-1}=o(1)$. To simplify the procedure and notations, here we just clarify the asymptotic properties of $N^{-1}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}(\widehat{Z}_{ij})$ without distinction of the null and alternatives. Our goal is to prove that
$\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\pi_{k}\left(\widehat{Z}_{ij}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left\{\pi_{k}(Z_{ij})-\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]e_{ij}-\frac{\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right]}{2}\left(e_{ij}^{2}-1\right)\right\}$
@@ A.4 Proof of Theorem 4, 5 and 6
We first focus on the estimation effects of $\widehat{\mu}_{j}-\mu_{j}$ and $\widehat{\sigma}^{2}-\sigma^{2}$. For $\widehat{\mu}_{j}-\mu_{j}$, by CLT, we have $\sqrt{N_{j}}(\widehat{\mu}_{j}-\mu_{j})\overset{d}{\to}\mathcal{N}(0,\sigma^{2})$ for each $1\leq j\leq J$ as $\min\{N_{1},\ldots,N_{J}\}\to\infty$. Besides, since
$\mathbb{E}\left[\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)^{2}\right]=\frac{\sigma^{2}}{N}\sum_{j=1}^{J}\mathbb{E}\left[N_{j}\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{2}\right]=\frac{J}{N}\sigma^{2},$ (A.16)
we have $N^{-1}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)^{2}=O_{p}(J/N)=o_{p}(N^{-1/2})$ from $J=o(N^{1/2})$. Analogously, by
$\mathbb{E}\left[\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{4}\right]=\frac{1}{N_{j}^{4}}\left[3N_{j}(N_{j}-1)+N_{j}\mathbb{E}\left[e^{4}\right]\right]=\frac{3}{N_{j}^{2}}+\frac{\mathbb{E}\left[e^{4}\right]-3}{N_{j}^{3}},$
@@
$\mathbb{E}\left[\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)^{4}\right]$
$=$ $\frac{\sigma^{4}}{N}\sum_{j=1}^{J}\mathbb{E}\left[N_{j}\left(\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}e_{ij}\right)^{4}\right]=\frac{\sigma^{4}}{N}\sum_{j=1}^{J}\left(\frac{3}{N_{j}}+\frac{\mathbb{E}\left[e^{4}\right]-3}{N_{j}^{2}}\right),$ (A.17)
which implies $N^{-1}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)^{4}=N^{-1}O_{p}(\sum_{j=1}^{J}N_{j}^{-1})=o_{p}(N^{-1})$ since $\sum_{j=1}^{J}N_{j}^{-1}=o_{p}(1)$. For $\widehat{\sigma}^{2}-\sigma^{2}$, we have
$\widehat{\sigma}^{2}-\sigma^{2}=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\left(Y_{ij}-\widehat{\mu}_{j}\right)^{2}-\sigma^{2}\right)$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}\left(\varepsilon_{ij}^{2}-\sigma^{2}\right)-\frac{1}{N}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)^{2}$
@@
$\frac{1}{N\sigma}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}$
$=$ $\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}+o_{p}\left(\frac{1}{\sqrt{N}}\right).$ (A.18)
@@
$\frac{1}{N\sigma}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}$
@@
$\frac{1}{N\sigma}\sum_{j=1}^{J}N_{j}\left(\widehat{\mu}_{j}-\mu_{j}\right)\left\{\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)\right\}-\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\frac{1}{N}\sum_{j=1}^{J}\sum_{i=1}^{N_{j}}e_{ij}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}=1}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}$
$=$ $\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i=1}^{N_{j}}\left(\dot{\pi}_{k}(Z_{ij})\phi\left(e_{ij}\right)-\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{ij}$
$+\frac{1}{N}\sum_{j=1}^{J}\frac{1}{N_{j}}\sum_{i_{1}=1}^{N_{j}}\sum_{i_{2}\neq i_{1}}^{N_{j}}\left(\dot{\pi}_{k}(Z_{i_{1}j})\phi\left(e_{i_{1}j}\right)-\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e_{i_{2}j}$
@@
$\mathbb{E}[A_{N1}]=\frac{J}{N}\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)e\right],$
@@
$\operatorname{Var}[A_{N1}]=\operatorname{Var}\left[\left(\dot{\pi}_{k}(Z)\phi\left(e\right)-\mathbb{E}\left[\dot{\pi}_{k}(Z)\phi\left(e\right)\right]\right)e\right]\frac{1}{N^{2}}\sum_{j=1}^{J}\frac{1}{N_{j}},$
then $A_{N1}=o_{p}(N^{-1/2})$ since $J=o(N^{1/2})$ as $N\to\infty$. For $A_{N2}$, we have $\mathbb{E}[A_{N2}]=0$, and